OpenStudy (anonymous):

i need help on these. 1) find remaining zeros for 7x^5+4x^4+14x^3+8x^2-441x-252 AND 2) 3x^4+13x^+17x^2+117x-90 WHEN the only zero given for Both is -3i someone please help i have bee on these for too long.

OpenStudy (helder_edwin):

if you know something about complex numbers then, if -3i is a zero (root) of the given polynomials, then 3i is also a root. this means that both polynomials can be exactly divided by \[ \Large (x+3i)(x-3i)=x^2-(3i)^2=x^2+9 \]

OpenStudy (anonymous):

divede them both by x^2+9, and get the remainging zeros from the quotient?

OpenStudy (helder_edwin):

it HAS to be zero, otherwise something is wrong with the polynomials you wrote down

OpenStudy (anonymous):

right, but the first one should be 5 zeros, and i only have -3i, and 3i

OpenStudy (helder_edwin):

yes i know but the quotient you get from the division MIGHT be easier to factor i guess

OpenStudy (anonymous):

3 and 3?

OpenStudy (anonymous):

i have this paper full of numbers and nothing is working, i might just give it a little break i dno

OpenStudy (helder_edwin):

i worked with the second polynomial i got this: \[ \Large 3x^4+13x^3+17x^2+117x-90=(x^2+9)(3x^2+13x-10) \] right?